A gas has a volume of 2.8 L at a temperature of 27°C. What temperature (°C) is needed to expand the volume to 15 L?


1
Expert's answer
2021-01-13T13:54:42-0500

Temperature should be in Kelvin for calculations: 27oC = 300 K.

According to the Charles' law,


"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


Therefore,

"T_2=\frac{T_1V_2}{V_1}=\frac{300K\times15L}{2.8L}=1607K=1334^\circ{C}" (1300oC to 2 significant digits)


Answer: 1300oC